M11-5 (Algo) Calculating Accounting Rate of Return, Payback Period [LO 11-1, 11-2]

Blue Marlin Company is considering the purchase of new equipment for its factory. It will cost $241,000 and have a $48,200 salvage value in five years. The annual net income from the equipment is expected to be $26,510, and depreciation is $38,560 per year.

  Calculate Blue Marlin’s accounting rate of return and payback period for the equipment. (Do not round intermediate calculations. Round your Payback Period to 2 decimal places.)


* Please explain how to get the ARR! I keep getting the answer wrong.
Answer #1








Annual rate of return=Annual net income/Equipment cost = 26510/241000= 11%








Payback period=241000/(26510+38560)= 3.70 years



















answered by: ANURANJAN SARSAM